Suppose there were many drivers w/ diverse WTP (MB = 100- Q where Q is the # beach goers) but constant private cost of $15 to get to the beach. Graph the Private Marginal Benefit, Private Marginal Cost and Social Cost Curves assuming each traveler causes $17 in external damage. Imposing a toll of $_____ would cause drivers to consume the efficient number of trips taken.

Suppose there were many drivers w/ diverse WTP (MB= 100- Q where Q is the # beach goers) but constant private cost of $15 to get to the beach. Disregard the MD=$17 above. What is the smallest value of MD which would cause all trips to be socially wasteful?
Answer #1

MB = 100-Q, PMC = 15,

SMC = 15+17 = 32

At eqm, impose toll equals to the Marginal damage caused at socially efficient eqm

So toll = $ 17

.________________________________

at social efficiency

PMB = SMC

100-Q = 32

Q*s = 68

☺ PMB, SMC, PMC PMB (68, 32) smic SMC PMC 50 100

________

all trips will be socially wasteful, if PMB <= SMC

100-Q <= 15+ MD

85-Q <= MD

smallest MD = $ 85